OpenStudy (anonymous):

Give the line and slope intercept: y=1/6x - 2

OpenStudy (anonymous):

is the question asking for x and y intercepts?

OpenStudy (anonymous):

Yes. Give the line and slope intercepts.

OpenStudy (anonymous):

since the equation is already in slope-intercept form, the y-intercept is -2... to find the x-intercept, set y=0 and solve for x: 0 = 1/6x -2 can you solve this ???

OpenStudy (anonymous):

is it ?\[y=\frac{1}{6}x-2\]

OpenStudy (anonymous):

y=mx+c m=slope and c is the intercept

OpenStudy (anonymous):

dpalnc will give it a try brb while i work it vidrantg yes

OpenStudy (anonymous):

I'm down to 0 = 1/6x - 2 2 + 2 2 = 1/6x